4 3 ND 1 7 1 2 3 4 7 To find the blue shaded area above, we would calculate: f(x)dx = area Where: b = f(x) = Area =
Answer #1

The region region R shaded iu figt The line yz f(x) is Porno Ky24x) 7x function passes through (4,0) and (0,4) Its equation i
Answer #2

SOLUTION :


            b

Area = ∫  f(x) dx

           a

where, 


a = 1 , b = 2 ,


f(x) = - x + 4  (since slope is - 1, and y-intercept is 4)


So, 


            

Area 


   2

= ∫  (- x + 4) dx 

   1

                             2

=  [- x^2 / 2 + 4x ]

                             1


= [ (- 2^2 / 2 + 4*2) - (- 1^2 / 2 + 4*1)]


= [ 6  - 7/2 ]


= 5 / 2


= 2.5   (ANSWER)



Check :


Graphically , Area = (3 + 2)/2 * 1 = 2.5 (same as above).

answered by: Tulsiram Garg